Events A and B are mutually exclusive. Suppose P(A)=0.19 and P(B)=0.06.
What is the probability that neither A nor B will occur?
Use two decimal digits.

Answer by rothauserc(4718)

The probability of events A or B occurring = 0.19 + 0.06 = 0.25
:
The probability that neither A nor B will occur is 1 - 0.25 = 0.75
: